### v1

@@ 5.1.2. Baselines
Models DAG (ours) TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN xPatch PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
NP
24 0.202 0.237 0.236 0.266 0.219 0.249 0.284 0.301 0.246 0.287 0.210 0.266 0.252 0.303 0.273 0.310 0.234 0.278 0.249 0.294 0.282 0.337
360 0.521 0.451 0.600 0.475 0.539 0.501 0.601 0.498 0.576 0.528 0.531 0.508 0.587 0.534 0.538 0.529 0.523 0.461 0.530 0.498 0.601 0.552
Avg 0.362 0.344 0.418 0.371 0.379 0.375 0.443 0.400 0.411 0.408 0.371 0.387 0.420 0.418 0.405 0.419 0.378 0.370 0.390 0.396 0.442 0.444
PJM
24 0.057 0.143 0.075 0.166 0.095 0.195 0.106 0.214 0.072 0.166 0.088 0.191 0.096 0.208 0.115 0.244 0.077 0.168 0.116 0.239 0.096 0.208
360 0.130 0.218 0.140 0.231 0.133 0.219 0.177 0.279 0.131 0.228 0.135 0.254 0.177 0.285 0.162 0.281 0.131 0.220 0.149 0.279 0.164 0.278
Avg 0.093 0.180 0.108 0.198 0.114 0.207 0.142 0.246 0.102 0.197 0.112 0.223 0.137 0.246 0.139 0.262 0.104 0.194 0.133 0.259 0.130 0.243
BE
24 0.361 0.229 0.392 0.253 0.426 0.272 0.426 0.285 0.432 0.272 0.391 0.259 0.471 0.339 0.451 0.319 0.411 0.256 0.452 0.326 0.451 0.310
360 0.485 0.330 0.512 0.327 0.482 0.310 0.571 0.364 0.597 0.436 0.568 0.416 0.646 0.487 0.645 0.495 0.579 0.410 0.702 0.538 0.600 0.442
Avg 0.423 0.279 0.452 0.290 0.454 0.291 0.498 0.325 0.515 0.354 0.479 0.337 0.559 0.413 0.548 0.407 0.495 0.333 0.577 0.432 0.526 0.376
FR
24 0.355 0.171 0.366 0.208 0.543 0.253 0.418 0.255 0.384 0.251 0.390 0.226 0.459 0.348 0.454 0.296 0.424 0.235 0.518 0.368 0.422 0.265
360 0.473 0.268 0.489 0.273 0.465 0.261 0.551 0.308 0.607 0.403 0.575 0.370 0.648 0.468 0.641 0.452 0.556 0.350 0.658 0.452 0.572 0.362
Avg 0.414 0.219 0.427 0.241 0.504 0.257 0.484 0.281 0.496 0.327 0.483 0.298 0.554 0.408 0.547 0.374 0.490 0.293 0.588 0.410 0.497 0.314
DE
24 0.277 0.322 0.339 0.362 0.380 0.383 0.367 0.383 0.376 0.378 0.387 0.396 0.394 0.407 0.399 0.412 0.399 0.394 0.413 0.412 0.423 0.426
360 0.462 0.418 0.610 0.474 0.599 0.509 0.630 0.511 0.589 0.482 0.583 0.507 0.551 0.474 0.547 0.479 0.550 0.472 0.589 0.498 0.573 0.487
Avg 0.370 0.370 0.475 0.418 0.489 0.446 0.499 0.447 0.482 0.430 0.485 0.452 0.473 0.441 0.473 0.445 0.474 0.433 0.501 0.455 0.498 0.457
Energy
24 0.079 0.215 0.122 0.273 0.093 0.235 0.103 0.248 0.117 0.283 0.241 0.418 0.138 0.306 0.135 0.298 0.092 0.240 0.239 0.390 0.085 0.230
360 0.169 0.320 0.204 0.357 0.167 0.331 0.202 0.355 0.288 0.452 0.237 0.385 0.328 0.472 0.302 0.464 0.204 0.367 0.214 0.363 0.263 0.416
Avg 0.124 0.267 0.163 0.315 0.130 0.283 0.153 0.302 0.203 0.367 0.239 0.402 0.233 0.389 0.218 0.381 0.148 0.303 0.226 0.377 0.174 0.323
Sdwpfm1
24 0.351 0.400 0.558 0.533 0.366 0.421 0.474 0.488 0.551 0.564 0.355 0.473 0.364 0.445 0.418 0.503 0.669 0.580 0.374 0.454 0.417 0.501
360 0.495 0.522 0.845 0.684 0.597 0.528 0.492 0.526 0.646 0.577 0.497 0.532 0.510 0.535 0.476 0.566 0.592 0.556 0.497 0.550 0.480 0.562
Avg 0.423 0.461 0.701 0.609 0.482 0.474 0.483 0.507 0.599 0.570 0.426 0.502 0.437 0.490 0.447 0.534 0.630 0.568 0.435 0.502 0.449 0.531
Sdwpfm2
24 0.372 0.414 0.627 0.570 0.411 0.458 0.461 0.492 0.445 0.452 0.477 0.536 0.394 0.462 0.474 0.538 0.508 0.478 0.418 0.492 0.459 0.530
360 0.583 0.556 0.978 0.736 0.541 0.519 0.511 0.540 0.584 0.528 0.589 0.611 0.587 0.563 0.520 0.589 0.523 0.513 0.602 0.603 0.589 0.630
Avg 0.477 0.485 0.803 0.653 0.476 0.488 0.486 0.516 0.514 0.490 0.533 0.573 0.491 0.512 0.497 0.564 0.516 0.495 0.510 0.547 0.524 0.580
Sdwpfh1
24 0.408 0.438 0.651 0.587 0.401 0.460 0.434 0.489 0.527 0.513 0.548 0.585 0.576 0.627 0.511 0.582 0.524 0.505 0.422 0.505 0.557 0.613
360 0.489 0.534 0.841 0.700 0.557 0.523 0.472 0.527 0.551 0.519 0.566 0.601 0.497 0.569 0.643 0.694 0.472 0.504 0.513 0.548 0.600 0.635
Avg 0.448 0.486 0.746 0.643 0.479 0.491 0.453 0.508 0.539 0.516 0.557 0.593 0.537 0.598 0.577 0.638 0.498 0.505 0.468 0.527 0.579 0.624
Sdwpfh2
24 0.438 0.465 0.820 0.677 0.474 0.493 0.579 0.553 0.629 0.563 0.468 0.540 0.473 0.533 0.580 0.614 0.640 0.554 0.457 0.527 0.633 0.653
360 0.608 0.595 0.962 0.761 0.657 0.549 0.619 0.614 0.665 0.569 0.608 0.609 0.569 0.628 0.713 0.729 0.579 0.547 0.641 0.632 0.704 0.674
Avg 0.523 0.530 0.891 0.719 0.566 0.521 0.599 0.583 0.647 0.566 0.538 0.574 0.521 0.581 0.647 0.672 0.610 0.551 0.549 0.580 0.669 0.664
Colbun
10 0.061 0.094 0.113 0.172 0.092 0.135 0.089 0.131 0.089 0.134 0.071 0.102 0.071 0.121 0.061 0.101 0.090 0.104 0.062 0.122 0.069 0.151
30 0.135 0.215 0.176 0.299 0.383 0.460 0.240 0.322 0.307 0.397 0.182 0.288 0.275 0.370 0.195 0.249 0.217 0.315 0.417 0.496 0.245 0.285
Avg 0.098 0.154 0.145 0.235 0.238 0.297 0.164 0.227 0.198 0.266 0.126 0.195 0.173 0.246 0.128 0.175 0.153 0.210 0.239 0.309 0.157 0.218
Rapel
10 0.151 0.203 0.301 0.308 0.201 0.253 0.228 0.271 0.174 0.219 0.163 0.209 0.181 0.227 0.174 0.231 0.161 0.216 0.163 0.218 0.184 0.239
30 0.309 0.408 0.387 0.416 0.409 0.414 0.411 0.432 0.365 0.432 0.340 0.417 0.333 0.416 0.325 0.390 0.505 0.546 0.374 0.445 0.316 0.410
Avg 0.230 0.305 0.344 0.362 0.305 0.333 0.320 0.351 0.269 0.326 0.252 0.313 0.257 0.321 0.249 0.311 0.333 0.381 0.269 0.332 0.250 0.324
$1^{st}$ Count 26 29 0 0 5 3 1 0 0 0 0 0 2 0 1 1 1 3 0 0 0 0
Table 3. Ablation studies for DAG.
@@
$\mathcal{F}_{\theta_{1}}$ + $\mathcal{G}_{\theta_{2}}$ +
Temporal Causality
0.504 0.509 0.189 0.261 0.855 0.585 0.240 0.389 0.447 0.436
$\mathcal{F}_{\theta_{3}}$ + $\mathcal{G}_{\theta_{4}}$ +
Channel Causality
0.392 0.459 0.170 0.255 0.473 0.438 0.350 0.467 0.346 0.405
@@
5.2. Main Results
Comprehensive forecasting results are presented in Table 1 and Table 2 to demonstrate the performance of DAG on the TSF-X task. We have the following observations:
1) Compared with forecasters of various structures, DAG achieves outstanding predictive performance. In terms of absolute metrics, Table 1 shows that DAG surpasses the second-best baseline, TimeXer, with a 10.8% reduction in MSE and a 6.4% reduction in MAE. Likewise, in Table 2, DAG achieves the highest number of first-place rankings—26 for MSE and 29 for MAE—demonstrating clear superiority over other methods such as TimeXer and TFT.
2) Methods that natively support future exogenous variables (e.g., TiDE, TFT, TimeXer) generally perform better than those that do not (e.g., PatchTST, DLinear), even when enhanced via MLP fusion. DAG, by explicitly modeling causal relationships while leveraging exogenous inputs, not only competes with these top baselines but surpasses them in nearly all metrics. This demonstrates the strength of incorporating causality-aware mechanisms in TSF-X tasks.
5.3. Model Analyses
5.3.1. Ablation Studies
We compare the full version of DAG with the following variants across four datasets. The Electricity has a lookback length of 96 and forecasts 720 steps, while the other three datasets have a lookback length of 720 steps and forecast 360 steps—see Table 3.
(a) $\mathcal{G}_{\theta_{2}}$: Uses only $X^{\text{endo}}$ to predict $Y^{\text{endo}}$.
(b) $\mathcal{G}_{\theta_{4}}$: Uses only $Y^{\text{exo}}$ to predict $Y^{\text{endo}}$.
(c) $\mathcal{G}_{\theta_{2}}$ + $\mathcal{G}_{\theta_{4}}$: Uses both $X^{\text{endo}}$ and $Y^{\text{exo}}$ to predict $Y^{\text{endo}}$.
Figure 4. Parameter sensitivity studies of main hyper-parameters in DAG.
(d) $\mathcal{F}_{\theta_{1}}$ + $\mathcal{G}_{\theta_{2}}$ + Temporal Causality: Predicts $Y^{\text{endo}}$ using $X^{\text{endo}}$ while incorporating temporal causality and the corresponding temporal causality loss derived from predicting $Y^{\text{exo}}$ using $X^{\text{exo}}$.
(e) $\mathcal{F}_{\theta_{3}}$ + $\mathcal{G}_{\theta_{4}}$ + Channel Causality: Predicts $Y^{\text{endo}}$ using $Y^{\text{exo}}$ while incorporating channel causality and the corresponding channel causality loss derived from predicting $X^{\text{endo}}$ using $X^{\text{exo}}$.
We make the following observations: 1) using only historical endogenous variables $\mathcal{G}_{\theta_{2}}$ or only future exogenous variables $\mathcal{G}_{\theta_{4}}$ results in suboptimal forecasting performance, indicating that relying on a single input source is insufficient for high-quality predictions. 2) Combining both $\mathcal{G}_{\theta_{2}}$ + $\mathcal{G}_{\theta_{4}}$ leads to a significant performance improvement, highlighting their complementary roles in modeling. 3) Furthermore, introducing temporal causality module ($\mathcal{F}_{\theta_{1}}$ + $\mathcal{G}_{\theta_{2}}$ + Temporal Causality) improves performance compared to using only historical endogenous variables $\mathcal{G}_{\theta_{2}}$; similarly, incorporating channel causality module ($\mathcal{F}_{\theta_{3}}$ + $\mathcal{G}_{\theta_{4}}$ + Channel Causality) outperforms the variant that uses only future exogenous variables $\mathcal{G}_{\theta_{4}}$. These results validate the effectiveness of both causality modeling mechanisms in DAG. 4) Ultimately, the full DAG model, which integrates both temporal and channel causality, achieves the best results, demonstrating the importance of modeling dual causality structures for TSF-X tasks.
Table 4. Experimental results under the setting where only historical exogenous variables are used, without future exogenous variables. The inputs are ($X^{\text{endo}}$ and $X^{\text{exo}}$). Avg means the average results from two forecasting horizons. Red: the best, Blue: the 2nd best.
Models DAG TimeXer CrossLinear DUET PatchTST
Datsets mse mae mse mae mse mae mse mae mse mae
10 0.070 0.098 0.084 0.146 0.068 0.101 0.073 0.106 0.083 0.125
30 0.180 0.223 0.181 0.293 0.190 0.289 0.221 0.290 0.217 0.316
Colbun
Avg 0.125 0.160 0.132 0.219 0.129 0.195 0.147 0.198 0.150 0.221
24 0.117 0.264 0.138 0.293 0.102 0.246 0.108 0.254 0.108 0.254
360 0.193 0.347 0.206 0.360 0.299 0.425 0.211 0.362 0.298 0.428
Energy
Avg 0.155 0.305 0.172 0.326 0.201 0.336 0.160 0.308 0.203 0.341
24 0.420 0.404 0.501 0.445 0.438 0.423 0.424 0.403 0.503 0.450
360 0.786 0.564 0.818 0.568 0.832 0.592 0.895 0.623 0.889 0.604
DE
Avg 0.603 0.484 0.659 0.507 0.635 0.508 0.660 0.513 0.696 0.527
24 0.072 0.162 0.096 0.191 0.096 0.198 0.082 0.178 0.106 0.212
360 0.182 0.259 0.186 0.267 0.197 0.283 0.198 0.273 0.189 0.273
PJM
Avg 0.127 0.211 0.141 0.229 0.147 0.241 0.140 0.226 0.148 0.243
24 0.226 0.255 0.270 0.292 0.246 0.282 0.262 0.271 0.283 0.296
360 0.609 0.505 0.609 0.474 0.655 0.505 0.626 0.494 0.630 0.505
NP
Avg 0.417 0.380 0.440 0.383 0.451 0.394 0.444 0.383 0.457 0.401
5.3.2. Parameter Sensitivity
We also conduct parameter sensitivity studies of DAG. We make the following observations: 1) Figure 4a and Figure 4b illustrate the performance variations of the DAG model under different settings of the fusion weight $\lambda_{1}$ and the causality weight $\lambda_{2}$. Overall, the model remains stable across a wide range of values, with the optimal range for both $\lambda_{1}$ and $\lambda_{2}$ typically falling between 0.3 and 0.7, indicating that moderate weighting is beneficial for improving performance. 2) Figure 4c shows that model dimension has a certain impact on performance, but DAG maintains good stability across different configurations. Smaller dimensions help reduce computational cost, while dimensions in the range of 64 to 256 generally achieve a good balance between prediction accuracy and model complexity. 3) Figure 4d explores the effect of patch length on model performance. Due to differences in temporal dependencies across datasets, the optimal patch length varies; however, a patch size between 8 and 32 tends to yield favorable results. It is worth noting that very small patch lengths may lead to increased computational complexity, while excessively large patch lengths may weaken the model’s ability to capture local dependencies.
5.3.3. Not Using Future Exogenous Variables
6. Conclusion

### v2

@@ 5 Experiments
Models DAG (ours) TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
NP 0.362 0.344 0.418 0.371 0.379 0.375 0.443 0.400 0.411 0.408 0.371 0.387 0.420 0.418 0.405 0.419 0.390 0.396 0.442 0.444
PJM 0.093 0.180 0.108 0.198 0.114 0.207 0.142 0.246 0.102 0.197 0.112 0.223 0.137 0.246 0.139 0.262 0.133 0.259 0.130 0.243
BE 0.423 0.279 0.452 0.290 0.454 0.291 0.498 0.325 0.515 0.354 0.479 0.337 0.559 0.413 0.548 0.407 0.577 0.432 0.526 0.376
FR 0.414 0.219 0.427 0.241 0.504 0.257 0.484 0.281 0.496 0.327 0.483 0.298 0.554 0.408 0.547 0.374 0.588 0.410 0.497 0.314
DE 0.370 0.370 0.475 0.418 0.489 0.446 0.499 0.447 0.482 0.430 0.485 0.452 0.473 0.441 0.473 0.445 0.501 0.455 0.498 0.457
Energy 0.124 0.267 0.163 0.315 0.130 0.283 0.153 0.302 0.203 0.367 0.239 0.402 0.233 0.389 0.218 0.381 0.226 0.377 0.174 0.323
Sdwpfm1 0.423 0.461 0.701 0.609 0.482 0.474 0.483 0.507 0.599 0.570 0.426 0.502 0.437 0.490 0.447 0.534 0.435 0.502 0.449 0.531
Sdwpfm2 0.477 0.485 0.803 0.653 0.476 0.488 0.486 0.516 0.514 0.490 0.533 0.573 0.491 0.512 0.497 0.564 0.510 0.547 0.524 0.580
Sdwpfh1 0.448 0.486 0.746 0.643 0.479 0.491 0.453 0.508 0.539 0.516 0.557 0.593 0.537 0.598 0.577 0.638 0.468 0.527 0.579 0.624
Sdwpfh2 0.523 0.530 0.891 0.719 0.566 0.521 0.599 0.583 0.647 0.566 0.538 0.574 0.521 0.581 0.647 0.672 0.549 0.580 0.669 0.664
Colbun 0.098 0.154 0.145 0.235 0.238 0.297 0.164 0.227 0.198 0.266 0.126 0.195 0.173 0.246 0.128 0.175 0.239 0.309 0.157 0.218
Rapel 0.230 0.305 0.344 0.362 0.305 0.333 0.320 0.351 0.269 0.326 0.252 0.313 0.257 0.321 0.249 0.311 0.269 0.332 0.250 0.324
1st Count 10 11 0 0 1 1 0 0 0 0 0 0 1 0 0 0 0 0 0 0
5.1 Experimental Settings
5.1.1 Datasets
To ensure comprehensive and fair comparisons across different models, we conduct experiments on 12 real-world datasets that satisfy the TSF-X conditions, including 5 EPF (Lago et al., 2021; Wang et al., 2024) datasets and 7 datasets we sourced ourselves. The future exogenous variables in these datasets are either approximately known or highly accurate. More details about the datasets are provided in Appendix A.1, Table 4.
5.1.2 Baselines
We comprehensively compare our model against 10 baselines, including 1) methods that inherently support future exogenous variables, such as TimeXer (Wang et al., 2024), TFT (Lim et al., 2021), and TiDE (Das et al., 2023), 2) as well as methods that do not originally support future exogenous variables, like DUET (Qiu et al., 2025), CrossLinear (Zhou et al., 2025), Amplifier (Fei et al., 2025), TimeKAN (Huang et al., 2025), xPatch (Stitsyuk and Choi, 2025), PatchTST (Nie et al., 2023), and DLinear (Zeng et al., 2023). For methods that do not support future covariates, we adapt them to incorporate future exogenous variables using an MLP fusion approach—see Algorithm 1 in Appendix A.3. Note that, for fairness, all methods in our experiments use historical endogenous variables, historical exogenous variables, and future exogenous variables.
5.1.3 Implementation Details
1) To keep consistent with previous works, we adopt Mean Squared Error (mse) and Mean Absolute Error (mae) as evaluation metrics. 2) We conduct both long-term and short-term prediction experiments. For the Colbun and Raperl datasets, the short-term prediction uses a lookback window of 60 with a prediction horizon of 10, while the long-term prediction uses a lookback window of 180 with a prediction horizon of 30. For the other datasets, the short-term prediction uses a lookback window of 168 with a prediction horizon of 24, and the long-term prediction uses a lookback window of 720 with a prediction horizon of 360. 3) We utilize the TFB (Qiu et al., 2024) code repository for unified evaluation, with all baseline results also derived from TFB. Following the settings in TFB (Qiu et al., 2024) and TSFM-Bench (Li et al., 2025), we do not apply the “Drop Last” trick to ensure a fair comparison. Further implementation details can be found in the Appendix A.2.
Table 2: Average results under the setting where only historical exogenous variables are used, without future exogenous variables. The inputs are ($X^{\text{endo}}$ and $X^{\text{exo}}$). Red: the best, Blue: the 2nd best. Full results are available in Table 6 in Appendix B.1.
Models DAG TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
NP 0.419 0.380 0.440 0.383 0.647 0.488 0.467 0.416 0.444 0.383 0.451 0.394 0.520 0.448 0.484 0.435 0.457 0.401 0.471 0.423
PJM 0.126 0.210 0.141 0.229 0.200 0.270 0.158 0.253 0.140 0.226 0.147 0.241 0.152 0.245 0.175 0.270 0.148 0.243 0.154 0.250
BE 0.462 0.297 0.477 0.301 0.563 0.354 0.547 0.348 0.473 0.305 0.477 0.300 0.502 0.325 0.488 0.315 0.485 0.309 0.549 0.381
FR 0.435 0.249 0.454 0.247 0.535 0.288 0.494 0.290 0.468 0.262 0.476 0.257 0.494 0.286 0.491 0.276 0.470 0.264 0.494 0.302
DE 0.603 0.487 0.659 0.507 0.684 0.515 0.644 0.519 0.660 0.513 0.635 0.508 0.712 0.548 0.669 0.526 0.696 0.527 0.690 0.529
Energy 0.150 0.300 0.172 0.326 0.376 0.480 0.162 0.311 0.160 0.308 0.201 0.336 0.180 0.327 0.147 0.296 0.203 0.341 0.160 0.307
Sdwpfm1 0.689 0.604 0.701 0.609 0.984 0.728 0.713 0.633 0.724 0.583 0.809 0.694 0.843 0.685 0.725 0.672 0.733 0.651 0.693 0.681
Sdwpfm2 0.786 0.657 0.803 0.653 1.044 0.767 0.816 0.682 0.820 0.641 0.800 0.687 0.942 0.732 0.836 0.701 0.834 0.714 0.805 0.728
Sdwpfh1 0.733 0.643 0.746 0.643 0.793 0.698 0.808 0.699 0.779 0.644 0.768 0.702 1.100 0.820 0.804 0.720 0.804 0.717 0.783 0.751
Sdwpfh2 0.870 0.695 0.891 0.719 0.926 0.746 0.919 0.751 1.007 0.715 0.956 0.774 0.980 0.768 0.941 0.782 1.025 0.802 0.913 0.798
Colbun 0.117 0.155 0.132 0.219 0.556 0.386 0.188 0.240 0.147 0.198 0.129 0.195 0.152 0.210 0.201 0.253 0.150 0.221 0.198 0.247
Rapel 0.261 0.281 0.344 0.362 0.308 0.334 0.323 0.357 0.304 0.306 0.240 0.299 0.337 0.348 0.342 0.337 0.325 0.334 0.319 0.338
1st Count 10 8 0 1 0 0 0 0 0 2 1 0 0 0 1 1 0 0 0 0
5.2 Main Results
Comprehensive forecasting results are presented in Table 1 to demonstrate the performance of DAG on the TSF-X task. We have the following observations:
1) Compared with forecasters of various structures, DAG achieves outstanding predictive performance. In terms of absolute metrics, Table 1 shows that DAG achieves the highest number of first-place rankings 10 for MSE and 11 for MAE demonstrating clear superiority over other methods such as TimeXer and TFT. 2) Methods that natively support future exogenous variables (e.g., TFT, TimeXer) generally perform better than those that do not (e.g., PatchTST, DLinear), even when enhanced via MLP fusion. DAG, by explicitly modeling correlation relationships while leveraging exogenous inputs, not only competes with these top baselines but surpasses them in nearly all metrics. This demonstrates the strength of incorporating correlation-aware mechanisms in TSF-X tasks.
5.3 Model Analyses
5.3.1 Not Using Future Exogenous Variables
5.3.2 Ablation Studies
We compare the full version of DAG with the following variants. The Electricity has a lookback length of 96 and forecasts 720 steps, while the other three datasets have a lookback length of 720 steps and forecast 360 steps—see Table 3. We make the following observations: 1) using only historical endogenous variables $\mathcal{G}_{\theta_{2}}$ or only future exogenous variables $\mathcal{G}_{\theta_{4}}$ results in suboptimal forecasting performance, indicating that relying on a single input source is insufficient for high-quality predictions. 2) Combining both $\mathcal{G}_{\theta_{2}}$ + $\mathcal{G}_{\theta_{4}}$ leads to a significant performance improvement, highlighting their complementary roles in modeling. 3) Furthermore, introducing temporal correlation module ($\mathcal{F}_{\theta_{1}}$ + $\mathcal{G}_{\theta_{2}}$ + Temporal Correlation) improves performance compared to using only historical endogenous variables $\mathcal{G}_{\theta_{2}}$; similarly, incorporating channel correlation module ($\mathcal{F}_{\theta_{3}}$ + $\mathcal{G}_{\theta_{4}}$ + Channel Correlation) outperforms the variant that uses only future exogenous variables $\mathcal{G}_{\theta_{4}}$. These results validate the effectiveness of both correlation modeling mechanisms in DAG. 4) Ultimately, the full DAG model, which integrates both temporal and channel correlation, achieves the best results, demonstrating the importance of modeling dual correlation structures for TSF-X tasks.
Table 3: Ablation studies for DAG.
@@ 5.3.2 Ablation Studies
(d) $\mathcal{F}_{\theta_{1}}$ + $\mathcal{G}_{\theta_{2}}$ + Temporal Correlation 0.504 0.509 0.189 0.261 0.855 0.585 0.240 0.389 0.447 0.436
(e) $\mathcal{F}_{\theta_{3}}$ + $\mathcal{G}_{\theta_{4}}$ + Channel Correlation 0.392 0.459 0.170 0.255 0.473 0.438 0.350 0.467 0.346 0.405
@@
6 Conclusion